# Function-Correcting Codes for Insertion-Deletion Channels

### v1

@@ IV VT-syndrome function
Definition 16 (VT-syndrome function).
@@
Using Theorem 2 and Corollary 1, we establish an upper bound and lower bound on $r^{VT}_{ID}(k,t)$ in the next lemma.
Lemma 14.
Let $f:\mathbb{F}_{2}^{k}\rightarrow\{0,1,\ldots,k\}$ be the VT-syndrome function defined by $f_{i}=i-1$ for all $1\leq i\leq k+1$, and let $x_{1}=(0,\ldots,0)$, $x_{i}=(0^{i-2}10^{k-i+1})$ denote the set of $k+1$ representative binary vectors of length $k$, where $2\leq i\leq k+1$. Consider the function distance matrix $I_{f}^{(2)}((t,f_{1},\ldots,f_{k+1});k)$ corresponding to $f$, and the insdel distance matrix $I_{f}^{(1)}(t,x_{1},\ldots,x_{k+1})$ corresponding to the set of representative binary vectors. Then,
$N_{f}^{(1)}(I_{\text{ID}}^{(1)}(t,x_{1},\ldots,x_{k+1}))\;\leq\;r^{\text{VT}}_{\text{ID}}(k,t)\;\leq\;N_{f}^{(2)}(I_{f}^{(2)}((t,f_{1},\ldots,f_{k+1});k)),$
@@
$[I_{f}^{(2)}((t,f_{1},\ldots,f_{k+1});k)]_{ij}=\begin{cases}0,&i=j,\\[4.0pt] 2(t+k),&i\neq j,\end{cases}$
@@
$I_{f}^{(2)}((t,f_{1},\ldots,f_{k+1});k)]_{ij}=\begin{cases}0,&i=j,\\[4.0pt] 2(t+k),&i\neq j,\end{cases}$
@@
Remark 2.
Since the non-diagonal entries of both the distance matrices $(I_{f}^{(1)}(t,x_{1},\ldots,x_{k+1})$ and $I_{\text{ID}}^{(2)}(t,x_{1},\ldots,x_{k+1})$ are equal to the same value $2t$ and $2(t+k)$ respectively, the bound on the optimal redundancy in Lemma 14 can be expressed as:
$N_{f}^{(1)}(k+1,2t)\;\leq\;r^{\text{VT}}_{\text{ID}}(k,t)\;\leq\;N_{f}^{(2)}(k+1,2(t+k);k))$
The following example illustrates Lemma 14 for particular values of $k$ and $t$.
Example 6.
Let $k=2$ and $t=1$. Then, from Lemma 14, the function insdel distance matrix corresponding to the VT-syndrome function and the insdel distance matrix corresponding to the binary vectors $x_{1}=(00)$, $x_{2}=(10)$, and $x_{3}=(01)$ are given respectively by
$I_{f}^{(2)}((1,f_{1},f_{2},f_{3});2)]_{ij}=\begin{pmatrix}0&6&6\\[4.0pt] 6&0&6\\[4.0pt] 6&6&0\end{pmatrix}$,
@@
$r^{\text{VT}}_{\text{ID}}(2,1)\leq N_{f}^{(2)}(I_{f}^{(2)}((1,f_{1},f_{2},f_{3});2))\leq 6$
V Number of Runs Function
@@
Definition 17.
A total number of runs function is defined as $f(x)=r(x)$, where $r(x)=|R(x)|$, $u\in\mathbb{F}^{k}_{2}$, and $k\in\mathbb{N}$.
Example 7.
@@
The expressiveness of this function is given by $E=|Im(r(\cdot))|=k$. Using Corollary 1 and Theorem 2, we establish bounds on the optimal redundancy of FCIDCs for the function $r(\cdot)$, as stated in the following lemma.
Lemma 15.
Let $f(x)=r(x)$ be the number of runs function on $x\in\mathbb{F}_{2}^{k}$. Consider the set of $k$ binary vectors defined as $x_{i}=0^{k-i+1}(10)^{i-1}/2$ if $i$ is odd and $x_{i}=0^{k-i+1}(10)^{(i-2)/2}1$ if $i$ is even. Then the optimal redundancy of FCIDC corresponding to $f$ is bounded as follows
$\scriptstyle N^{(1)}_{ID}(I_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k}))\leq r_{ID}^{f}(k,t)\leq N^{(2)}_{ID}(I_{f}^{(2)}(t,f_{1},f_{2},\dots,f_{k});k)$
Where $I_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k})$ and $I_{f}^{(2)}(t,f_{1},f_{2},\dots,f_{k});k)$ are order $k$ distance matrix and function distance matrix, respectively, and their entries are as follows:
@@
Therefore, the entries of the function insdel matrix and insdel distance matrices associated with the set $\{x_{i}\}_{i=1}^{k+1}$ are given respectively by
@@
Example 8.
Let $k=4$ and $t=1$. Then, from Lemma 15, the function distance matrix corresponding to the number of runs function and the distance matrix corresponding to the binary vectors $x_{1}=(0000)$, $x_{2}=(0001)$, $x_{3}=(0010)$ and $x_{4}=(0101)$ are given respectively by
$I_{f}^{(2)}((1,f_{1},f_{2},f_{3});2)]_{ij}=\begin{pmatrix}0&10&10&8\\[4.0pt] 10&0&10&10\\[4.0pt] 10&10&0&10\\[4.0pt] 8&10&10&0\end{pmatrix}$,
@@
The next result presents a lower bound on the redundancy of the number of runs function that is inferred from Lemma 15 and Plotkin-like bound of Lemma 11
Corollary 4.
@@
[…]
@@
For $1\leq i\leq k$, let $x_{i}=0^{k-i+1(10)^{i-1}/2}$ if $i$ is odd and $x_{i}=0^{k-i+1}(10)^{(i-2)/2}1$ if $i$ is even. Let $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{k}\}$ be a $I_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k})$-code of length $N^{(1)}_{ID}(I_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k}))$. Consider the first $t+2$ codewords of $\mathcal{P}$ and $(t+2)\times(t+2)$ leading principal submatrix of $I_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k})$ denoted by $I_{t+2}.$ Then, $\{p_{1},p_{2},\ldots,p_{t+2}\}$ is a $I_{t+2}$-code. By Corollary 3,
@@
From Lemma 15, we have
$r_{ID}^{f}(k,t)$ $\geq N^{(1)}_{ID}(I_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k}))$
@@
Next, we present a construction for FCIDCs for the run-length function which utilises the concept of the shifted modulo operator defined in [10132545]. The shifted modulo operator is defined as follows:
Definition 18.
@@
$Enc(x)=(x,p_{r(x)}),$
where the $p_{i}$’s are defined depending on $t$ as follows. For $t=1$, set $p_{1}=(00),p_{2}=(10),$ and $p_{3}=(01)$. Set $p_{i}=p_{i\text{ }smod\text{ }3}$ for $i\geq 4$.
For $t=2$, set $p_{1}=(00000),p_{2}=(00011),p_{3}=(01010),p_{4}=(01111)$ and $p_{5}=11000$. Set $p_{i}=p_{i\text{ }smod\text{ }3}$ for $i\geq 6$.
for $t\geq 3$, let $p_{1},\ldots,p_{2t+1}$ be a code with minimum distance $2t$, i.e., $d_{ID}(p_{i},p_{j})\geq 2t\quad\forall i,j\leq 2t+1,i\neq j$ and set $p_{i}=p_{i\text{ }smod\text{ }(2t+1)}$ for $i\geq 2t+2$.
VI Maximum Run-Length Function
@@
Definition 19 (Maximum Run-Length Function).
@@
Let $y^{\prime}$ be a substring of $y$ aligned with the run $0^{i}$. If $t$ is the number of $1s$ in $y^{\prime}$, then the number of zeroes in $y^{\prime}$ is iti-t. This means $y^{\prime}$ can have at most $t+1$ blocks of zeroes, each split by $1$ and having size at most $j$. Hence, we have,
@@
The following example shows that the bound in Lemma 16 is tight.
Example 10.
@@
Using Lemma 16 and Corollary 2, we give an upper bound on the optimal redundancy of FCIDCs designed for maximum run-length function.
Lemma 17.
@@
By Lemma 16, $d_{ID}(x,y)\leq 2\Big\lceil\frac{|i-j|}{min(i,j)+1}\Big\rceil\quad\forall 1\leq i,j\leq k.$ Let $\bm{I}$ be a square matrix of order $k$ whose entries are:
@@
Then, by Corollary 2 $r_{ID}^{f}(k,t)\leq N_{ID}^{(2)}(\bm{I};k).$
The next lemma gives a lower bound on the optimal redundancy of FCIDCs designed for the maximum run-length functions by using Corollary 1.
Lemma 18.
@@
From Lemma 2 we have $d_{ID}(x_{i},x_{j})\geq|r(x_{i})-r(x_{j})|=|(1+k-i)-(1+k-j)|=|j-i|=i-j.$
@@
From Lemma 2 we have $d_{ID}(x_{i},x_{j})\geq|r(x_{i})-r(x_{j})|+1=|(1+k-i)-(1+k-j)|+1=|j-i|+1=i-j+1.$
@@
and the lower bound on the redundancy, $r_{ID}^{f}(k,t)\geq N_{ID}^{(1)}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\ldots,x_{k}))$ follows from Corollary 1.
Example 11.
Let $k=4$ and $t=1$. Then, from Lemma 15, the function distance matrix corresponding to the number of runs function and the distance matrix corresponding to the binary vectors $x_{1}=(0000)$, $x_{2}=(0001)$, $x_{3}=(0010)$ and $x_{4}=(0101)$ are given respectively by
$I_{f}^{(2)}((1,f_{1},f_{2},f_{3});2)]_{ij}=\begin{pmatrix}0&10&10&8\\[4.0pt] 10&0&10&10\\[4.0pt] 10&10&0&10\\[4.0pt] 8&10&10&0\end{pmatrix}$,
@@
VII Locally $(\lambda,\rho)_{I}$-function
Lenz et al. first studied a class of functions, called locally binary[10132545] with respect to the Hamming metric and derived optimal function-correcting codes (FCCs) for them. This concept was subsequently generalized by Rajput et al. to locally $(\lambda,\rho)$-functions[Rajput2025] for the Hamming metric and later by Verma et al. for the $b$-symbol metric [11250740]. The significance of this class of functions stems from the fact that any function with a finite image set can be represented as a locally $(\lambda,\rho)$-function. This universality is the primary reason the theory, initially developed for the Hamming metric, has been extended to other metric spaces where FCCs are being explored. In this section, we study the class of locally $(\lambda,\rho)$-functions in the insdel metric setting.
Definition 20 (Function ball).
The function ball of a function $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ with radius $\rho$ around $u$ in insdel metric is defined as
$B_{ID}^{f}(u,\rho)=f(B_{ID}(u,\rho))=\{f(u)|u\in B_{ID}(u,\rho)\}.$
Definition 21 (Locally bounded function in insdel metric).
A locally $(\rho,\lambda)_{ID}$ function $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ is a function for which $|B_{ID}^{f}(u,\rho)|\leq\lambda,\quad\forall u\in\mathbb{F}_{2}^{k}.$
The following lemma is a straightforward extension of Lemma 6 to the insdel metric setting. This result has subsequently been employed to derive an upper bound on the optimal redundancy of FCIDC corresponding to locally $(\rho,\lambda)_{ID}$-functions.
Lemma 19.
For any locally $(\rho,\lambda)_{ID}$-function $f:\mathbb{F}_{2}^{k}\to Im(f)$, there exists a mapping $\mathrm{Col}_{f}:\mathbb{F}_{2}^{k}\to[\lambda]$ such that for any $u,v\in\mathbb{F}_{2}^{k}$ with $f(u)\neq f(v)$ and $d_{ID}(u,v)\leq\rho$, we have $\mathrm{Col}_{f}(u)\neq\mathrm{Col}_{f}(v)$.
We now present an upper bound on the optimal redundancy of FCIDC corresponding to locally $(\rho,\lambda)_{ID}$-functions, expressed in terms of the shortest possible length of a binary insdel code with a prescribed number of codewords and minimum insdel distance.
Theorem 3.
Let $t$ be a positive integer and $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ be a locally $(\lambda,2t)_{ID}$-function then $r_{ID}^{f}(k,t)\leq N^{(2)}_{ID}(\lambda,2(t+k);k)$.

### v2

@@ 4 VT-syndrome function
Definition 4.1 (VT-syndrome function).
@@
Using Theorem 3.22 and Corollary 3.19, we establish an upper bound and lower bound on $r^{VT}_{ID}(k,t)$ in the next lemma.
Lemma 4.2.
Let $f:\mathbb{F}_{2}^{k}\rightarrow\{0,1,\ldots,k\}$ be the VT-syndrome function defined by $f_{i}=i-1$ for all $1\leq i\leq k+1$, and let $x_{1}=(0,\ldots,0)$, $x_{i}=(0^{i-2}10^{k-i+1})$ denote the set of $k+1$ representative binary vectors of length $k$, where $2\leq i\leq k+1$. Consider the function distance matrix $\bm{I}_{f}^{(2)}((t,f_{1},\ldots,f_{k+1});k)$ corresponding to $f$, and the distance matrix $I_{f}^{(1)}(t,x_{1},\ldots,x_{k+1})$ corresponding to the set of representative binary vectors. Then,
$N_{ID}^{(1)}(I_{\text{ID}}^{(1)}(t,x_{1},\ldots,x_{k+1}))\;\leq\;r^{\text{VT}}_{\text{ID}}(k,t)\;\leq\;N_{ID}^{(2)}(\bm{I}_{f}^{(2)}((t,f_{1},\ldots,f_{k+1});k)),$
@@
$[\bm{I}_{f}^{(2)}((t,f_{1},\ldots,f_{k+1});k)]_{ij}=\begin{cases}0,&i=j,\\[4.0pt] 2(t+k),&i\neq j,\end{cases}$
@@
$\bm{I}_{f}^{(2)}((t,f_{1},\ldots,f_{k+1});k)]_{ij}=\begin{cases}0,&i=j,\\[4.0pt] 2(t+k),&i\neq j,\end{cases}$
@@
Remark 4.3.
Since the non-diagonal entries of both the distance matrices $(I_{f}^{(1)}(t,x_{1},\ldots,x_{k+1})$ and $I_{\text{ID}}^{(2)}(t,x_{1},\ldots,x_{k+1})$ are equal to the same value $2t$ and $2(t+k)$ respectively, the bound on the optimal redundancy in Lemma 4.2 can be expressed as:
$N_{ID}^{(1)}(k+1,2t)\;\leq\;r^{\text{VT}}_{\text{ID}}(k,t)\;\leq\;N_{ID}^{(2)}(k+1,2(t+k);k))$
The next corollary gives an explicit bound on the optimal redundancy of FCIDCs for the VT syndrome function. The upper bound can be derived by using the result of the preceding remark together with Theorem 3.30, whereas the lower bound follows from the Remark 4.3 and Lemma 3.28.
Corollary 4.4.
Let $k\geq 2$ and $t\in\mathbb{N}$, then the optimal redundancy of FCIDCs for the VT syndrome function defined over $\mathbb{F}_{2}^{k}$ is bounded above as follows:
$\frac{2t(k+1)}{k+2}\leq r^{\text{VT}}_{\text{ID}}(k,t)\leq\Bigg\lceil\frac{\ln(k+1)+2(t+k-1)\ln\Big(\frac{ek}{t+k-1}\Big)}{\ln 2}\Bigg\rceil.$
The following example illustrates Lemma 4.2 for particular values of $k$ and $t$.
Example 4.5.
Let $k=2$ and $t=1$. Then, from Lemma 4.2, the function distance matrix corresponding to the VT-syndrome function and the distance matrix corresponding to the binary vectors $x_{1}=(00)$, $x_{2}=(10)$, and $x_{3}=(01)$ are given respectively by
$\bm{I}_{f}^{(2)}((1,f_{1},f_{2},f_{3});2)]_{ij}=\begin{pmatrix}0&6&6\\[4.0pt] 6&0&6\\[4.0pt] 6&6&0\end{pmatrix}$,
@@
$r^{\text{VT}}_{\text{ID}}(2,1)\leq N_{f}^{(2)}(\bm{I}_{f}^{(2)}((1,f_{1},f_{2},f_{3});2))\leq 6$
5 Number of Runs Function
@@
Definition 5.1.
A total number of runs function is defined as $f(x)=r(x)$, where $r(x)=|R(x)|$, $x\in\mathbb{F}^{k}_{2}$, and $k\in\mathbb{N}$.
Example 5.2.
@@
The expressiveness of this function is given by $E=|Im(r(\cdot))|=k$. Using Corollary 3.19 and Theorem 3.22, we establish bounds on the optimal redundancy of FCIDCs for the function $r(\cdot)$, as stated in the following lemma.
Lemma 5.3.
Let $f(x)=r(x)$ be the number of runs function on $x\in\mathbb{F}_{2}^{k}$. Consider the set of $k$ binary vectors defined as $x_{i}=0^{k-i+1}(10)^{(i-1)/2}$ if $i$ is odd and $x_{i}=0^{k-i+1}(10)^{(i-2)/2}1$ if $i$ is even. Then the optimal redundancy of FCIDCs corresponding to $f$ is bounded as follows
$N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k}))\leq r_{ID}^{f}(k,t)\leq N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,f_{1},f_{2},\dots,f_{k});k)$
Where $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k})$ and $\bm{I}_{f}^{(2)}(t,f_{1},f_{2},\dots,f_{k});k)$ are order $k$ distance matrix and function distance matrix, respectively, and their entries are as follows:
@@
Therefore, the entries of the function insdel matrix and insdel distance matrices associated with the set $\{x_{i}\}_{i=1}^{k}$ are given respectively by
@@
This establishes the desired bounds on $r^{f}_{\text{ID}}(k,t)$, completing the proof. ∎
Example 5.4.
Let $k=4$ and $t=1$. Then, from Lemma 5.3, the function distance matrix corresponding to the number of runs function and the distance matrix corresponding to the binary vectors $x_{1}=(0000)$, $x_{2}=(0001)$, $x_{3}=(0010)$ and $x_{4}=(0101)$ are given respectively by
$\bm{I}_{f}^{(2)}((1,f_{1},f_{2},f_{3});2)]_{ij}=\begin{pmatrix}0&10&10&8\\[4.0pt] 10&0&10&10\\[4.0pt] 10&10&0&10\\[4.0pt] 8&10&10&0\end{pmatrix}$,
@@
The next result presents a lower bound on the redundancy of the number of runs function that is inferred from Lemma 5.3 and Plotkin-like bound of Lemma 3.26
Corollary 5.5.
@@
$r_{\text{ID}}^{f}(k,t)\geq\frac{10t^{3}+39t^{2}+26t}{3(t+2)^{2}}\quad,\text{ when $t$ is even }$
and
$r_{\text{ID}}^{f}(k,t)\geq\frac{10t^{3}+39t^{2}+38t+29}{3(t+2)^{2}},\quad\text{ when $t$ is odd. }$
@@
For $1\leq i\leq k$, let $x_{i}=0^{k-i+1(10)^{(i-1)/2}}$ if $i$ is odd and $x_{i}=0^{k-i+1}(10)^{(i-2)/2}1$ if $i$ is even. Let $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{k}\}$ be a $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k})$-code of length $N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k}))$. Consider the first $t+2$ codewords of $\mathcal{P}$ and $(t+2)\times(t+2)$ leading principal submatrix of $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k})$ denoted by $I_{t+2}.$ Then, $\{p_{1},p_{2},\ldots,p_{t+2}\}$ is a $I_{t+2}$-code. By Corollary 3.27,
@@
From Lemma 5.3, we have
$r_{ID}^{f}(k,t)$ $\geq N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k}))$
@@
Next, we present a construction for FCIDCs for the number of runs function which utilises the concept of the shifted modulo operator defined in [1].
The shifted modulo operator is defined as follows:
Definition 5.6.
@@
$\psi(x)=(x,p_{r(x)}),$
where the $p_{i}$’s are defined depending on $t$ as follows.
For $t=1$, set $p_{1}=(00)^{k+1},\text{ }p_{2}=(10)^{k+1},$ and $p_{3}=(0^{k+1}1^{k+1})$. Set $p_{i}=p_{i\text{ }smod\text{ }3}$ for $i\geq 4$.
For $t=2$, set $p_{1}=(00)^{k+2},\text{ }p_{2}=(10)^{k+2},$ $p_{3}=(0^{k+2}1^{k+2})$ and $p_{4}=(1^{k+2}0^{k+2})$. Set $p_{i}=p_{i\text{ }smod\text{ }5}$ for $i\geq 5$.
For $t\geq 3$, let $p_{1},\ldots,p_{2t+1}$ be an insdel code with minimum insdel distance $2(k+t)$, i.e., $d_{ID}(p_{i},p_{j})\geq 2(t+k)\quad\forall i,j\leq 2t+1,\text{ }i\neq j$ and set $p_{i}=p_{i\text{ }smod\text{ }(2t+1)}$ for $i\geq 2t+2$.
The FCIDC construction gives an upper bound on the optimal redundancy of the number of runs function as stated in the following lemma.
Lemma 5.7.
Let $k,t\in\mathbb{N}$ and $f$ be the number of runs function defined on $\mathbb{F}_{2}^{k}$. Then, the optimal redundancy of FCIDCs corresponding to f is upper bounded as follows:
  • $r_{ID}^{f}(k,1)\leq 2(k+1)$.
  • $r_{ID}^{f}(k,2)\leq 2(k+2)$.
  • $r_{ID}^{f}(k,t)\leq N_{ID}^{(2)}(2t+1,2(t+k);k)\leq\Bigg\lceil\frac{\ln(2t+1)+2(t+k-1)\ln\Big(\frac{ek}{t+k-1}\Big)}{\ln 2}\Bigg\rceil$ .
The proof of this lemma follows from the construction given in the Appendix 9.
6 Maximum Run-Length Function
@@
Definition 6.1 (Maximum Run-Length Function).
@@
Let $y^{\prime}$ be a substring of $y$ aligned with the run $0^{i}$. If $t$ is the number of $1s$ in $y^{\prime}$, then the number of zeros in $y^{\prime}$ is iti-t. This means $y^{\prime}$ can have at most $t+1$ blocks of zeros, each split by $1$ and having size at most $j$. Hence, we have,
@@
The following example shows that the bound in Lemma 6.3 is tight.
Example 6.5.
@@
Using Lemma 6.3 and Corollary 3.24, we give an upper bound on the optimal redundancy of FCIDCs designed for maximum run-length function.
Lemma 6.6.
@@
By Lemma 6.3, $d_{ID}(x,y)\leq 2\Big\lceil\frac{|i-j|}{min(i,j)+1}\Big\rceil\quad\forall 1\leq i,j\leq k.$ Let $\bm{I}$ be a square matrix of order $k$ whose entries are:
@@
Then, by Corollary 3.24 $r_{ID}^{f}(k,t)\leq N_{ID}^{(2)}(\bm{I};k).$
Observe that $max_{i,j}[\mathbf{I}]_{ij}=2(t+k)$ in Lemma 6.6, therefore the following redundancy bound follows from Corollary 3.31.
Corollary 6.7.
Let $k,t\in\mathbb{N}$ and $r_{ID}^{f}(k,t)$ be the optimal redundancy of FCIDCs corresponding to $f(x)=r_{max}(x)$. Then,
$N_{ID}^{(2)}(k,2(t+k);k)\leq\Bigg\lceil\frac{\ln(k)+2(t+k-1)\ln\Big(\frac{ek}{t+k-1}\Big)}{\ln 2}\Bigg\rceil.$
The next lemma gives a lower bound on the optimal redundancy of FCIDCs designed for the maximum run-length functions by using Corollary 3.19.
Lemma 6.8.
@@
From Lemma 2.9 we have $d_{ID}(x_{i},x_{j})\geq|r(x_{i})-r(x_{j})|=|(1+k-i)-(1+k-j)|=|j-i|=i-j.$
@@
From Lemma 2.9 we have $d_{ID}(x_{i},x_{j})\geq|r(x_{i})-r(x_{j})|+1=|(1+k-i)-(1+k-j)|+1=|j-i|+1=i-j+1.$
@@
and the lower bound on the redundancy, $r_{ID}^{f}(k,t)\geq N_{ID}^{(1)}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\ldots,x_{k}))$ follows from Corollary 3.19.
Because the matrix entries $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\ldots,x_{k})$(Lemma 6.8) associated with maximum run length function match those for the number of runs (Lemma 5.3), the lower bound on optimal redundancy for FCIDCs is identical in both cases(Corollary 5.5), as given in the corollary below.
Corollary 6.9.
For any $k\geq t+2$. Then,
This simplifies to:
[…]
and
[…]
Example 6.10.
Let $k=4$ and $t=1$. Then, from Lemma 5.3, the function distance matrix corresponding to the number of runs function and the distance matrix corresponding to the binary vectors $x_{1}=(0000)$, $x_{2}=(0001)$, $x_{3}=(0010)$ and $x_{4}=(0101)$ are given respectively by
$\bm{I}_{f}^{(2)}((1,f_{1},f_{2},f_{3});2)]_{ij}=\begin{pmatrix}0&10&10&8\\[4.0pt] 10&0&10&10\\[4.0pt] 10&10&0&10\\[4.0pt] 8&10&10&0\end{pmatrix}$,
@@
7 Locally $(\lambda,\rho)_{ID}$-function
Lenz et al. first studied a class of functions, called locally binary function[1] with respect to the Hamming metric and derived optimal function-correcting codes (FCCs) for them. This concept was subsequently generalized by Rajput et al. to locally $(\lambda,\rho)$-functions[16] for the Hamming metric, and later by Verma et al. for the $b$-symbol metric [18]. The significance of this class of functions stems from the fact that any function with a finite image set can be represented as a locally $(\lambda,\rho)$-function. This universality is the primary reason the theory, initially developed for the Hamming metric, has been extended to other distance metrics where FCCs are being explored. In this section, we study the class of locally $(\lambda,\rho)_{ID}$-functions in the insdel metric setting.
Definition 7.1 (Function ball).
The function ball of a function $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ with radius $\rho$ around $u\in\mathbb{F}_{2}^{k}$ in insdel metric is defined as
$B_{ID}^{f}(u,\rho)=f(B_{ID}(u,\rho))=\{f(v)|v\in B_{ID}(u,\rho)\}.$
Definition 7.2 (Locally bounded function in insdel metric).
A locally $(\lambda,\rho)_{ID}$ function $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ is a function for which $|B_{ID}^{f}(u,\rho)|\leq\lambda,\quad\forall u\in\mathbb{F}_{2}^{k}.$
The following lemma is a straightforward extension of Lemma 2.17 to the insdel metric setting. This result has subsequently been employed to derive an upper bound on the optimal redundancy of FCIDCs corresponding to locally $(\lambda,\rho)_{ID}$-functions.
Lemma 7.3.
Let $f:\mathbb{F}_{2}^{k}\rightarrow\mathrm{Im}(f)$ be a locally $(\lambda,\rho)_{ID}$-function. Assume that $\mathrm{Im}(f)$ is equipped with a total order $\prec$, and that for every $u\in\mathbb{F}_{2}^{k}$, the set $B_{ID}^{f}(u,\rho)$ forms a contiguous block with respect to $\prec$. Then there exists a mapping
$\mathrm{Col}_{f}:\mathbb{F}_{2}^{k}\rightarrow[\lambda]$
such that for all $u,v\in\mathbb{F}_{2}^{k}$ satisfying $f(u)\neq f(v)\quad\text{and}\quad d_{ID}(u,v)\leq\rho$, we have $\mathrm{Col}_{f}(u)\neq\mathrm{Col}_{f}(v)$.
Throughout the remainder of this section, we assume that all locally $(\lambda,\rho)_{ID}$ functions under consideration satisfy the assumption in Lemma 7.3. We now present an upper bound on the optimal redundancy of FCIDCs corresponding to locally $(\lambda,\rho)_{ID}$-functions, expressed in terms of the shortest possible length of a binary insdel code with a prescribed number of codewords and minimum insdel distance.
Theorem 7.4.
Let $t$ be a positive integer and $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ be a locally $(\lambda,2t)_{ID}$-function then $r_{ID}^{f}(k,t)\leq N^{(2)}_{ID}(\lambda,2(t+k);k)\leq\frac{\ln\lambda+2(t+k-1)\ln\frac{ek}{t+k-1}}{\ln 2}$.